OpenStudy (anonymous):

how do i solve this 5(x-6)/7=3x+4 **the 5(x-6) is over 7**

OpenStudy (anonymous):

I'd start by multiplying both sides by 7: 5(x-6)=21x+28

OpenStudy (anonymous):

Then distribute that 5 on the left side: 5x-30=21x+28

OpenStudy (anonymous):

And from here, we can just combine terms: 5x-2=21x

OpenStudy (anonymous):

and even further, -2=16x

OpenStudy (anonymous):

So now you just divide both sides by 16, and you get 1/8 = x

OpenStudy (anonymous):

Sorry, -1/8 = x

OpenStudy (jwheele1):

5(x-6) ----- = 3x +4 7 /\/\/\/\/\/\/\//\/\/\/\/\/\//\/\/\/ Distribute: 5x-30 ----- = 3x +4 7 /\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/ make both sides fraction: 5x-30 3x +4 ----- = --------- 7 1 /\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/\/ cross multiply 1(5x-30) = 7(3x+4) /\/\/\/\/\/\/\/\/\/\/\/\/\/\/ distribute 5x-30 = 21x+28 /\/\/\/\/\/\/\/\/\/\/\/\ make it to where x is only on one side of the equation by subtracting 5x from both sides. 5x-30 = 21x+28 -5x -5x -30 = 16x +28 /\/\/\/\/\/\/\/\/\/\/\/\/ isolate the number glued to the variable (16x) by subtracting 28 from both sides. -30 = 16x +28 -28 -28 -58 = 16x /\/\/\/\/\/\/\/\/\/ divide both sides by the number glued to the variable (16) -29 ---- = x 8 ...this doesn't seem like a decent answer....perhaps I had an error?

OpenStudy (jwheele1):

Lets see koby's way instead: 5(x-6) ----- = 3x +4 7 /\/\/\/\/ 5(x-6) 7/1 { ----- } = 7(3x +4) 7 /\/\/\/\ on the left of the equal the 7 cancels out... 5(x-6) = 7(3x +4) /\/\/\/\/\/\/ distribute 5x-30 = 21x+28 /\////\/\/\/\/\/\/\/ lets add 30 to both sides. 5x-30 = 21x+28 +30 -30 5x =21x -2 /\/\/\/\/\/\/\/ then subtract 21x from both sides 5x =21x -2 -21x -21x -16x = -2 /\/\/\/\/\ divide both sides by -16 -16x = -2 ---- = ---- -16 -16 x = 1/8